OpenStudy (anonymous):

Solve x y 2x+3y=11 3x+y=6

jhonyy9 (jhonyy9):

y=6-3x 2x+3(6-3x)=2x+18-9x=18-7x=11 -7x=-7 x=1 3(1)+y=6 3+y=6 y=3 x=1 y=3